# ExDDV: A New Dataset for Explainable Deepfake Detection in Video

### v1

@@ 6 Conclusion
Acknowledgments. This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-2-0227, within PNCDI IV.
Supplementary Material
@@
7 Additional results
In this section, we present some additional data and more results. In Figure 7, we present more annotated samples from ExDDV, having different levels of difficulty, click locations, and explanatory text lengths.
Figure 8 contains a comprehensive diagram with qualitative samples for all possible training scenarios applied on LLaVA [45]. We observe that most explanations are meaningful, although there are cases when various model versions fail.
In Figure 10, we showcase a comparison between the predictions of the click regressor and the ground-truth click locations, highlighting the precise localization performance of our model.
@@ 8 Training environments
We have worked on multiple environments for our experiments. For the in-context learning experiments, we used a Tesla V100-SXM2 GPU with 32GB VRAM. Phi-3-Vision [1] and LLaVA [45] were fine-tuned on a single H100 GPU with 80GB VRAM. BLIP-2 [40] was fine-tuned using an RTX 4090 GPU with 24GB VRAM. The same training environment as for BLIP-2 was used for the click predictors.
Figure 9: The number of videos for each frame resolution in ExDDV.
Figure 10: Qualitative examples of our click predictor. The green dots represent the ground-truth locations and the blue dots the corresponding predictions. Best viewed in color.

### v2

@@ 6 Conclusion
Acknowledgments
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-2-0227, within PNCDI IV. This research was also supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 351416.
7 Annotation Process and Video Resolutions
The GUI (shown in Fig. 2) contains two video players, side by side. The deepfake video is played on the left side of the window, while the corresponding real video is played on the right side. By default, the real video is not played. If the annotator needs to play the real video along with the deepfake one, they could simply press the button below the player screen. When the fake video is playing, the user can click anywhere on the frame to indicate the location of artifacts. Behind the scene, the application records the relative pixel location and the timestamp (i.e. frame index) of the click. Under the fake video player, there is a text box in which the details describing the visual issues can be written by the user. In the bottom right, there are three radio buttons, which allow the user to indicate the difficulty level of identifying deepfake evidence. We ask users to label deepfake videos as hard, when they need to play the deepfake video at least two times, or when they need to activate the real video to observe artifacts. In a similar manner, we instruct them to label videos as easy, if they are able to identify more than one artifact with a single play of the deepfake video. For real videos, we do not collect clicks.
In Figure 9, we present more annotated samples from ExDDV, having different levels of difficulty, click locations, and explanatory text lengths.
Figure 8: The number of videos for each frame resolution in ExDDV.
@@ 7 Annotation Process and Video Resolutions
8 Additional Results
Figure 10 contains a comprehensive diagram with qualitative samples for all possible training scenarios applied on LLaVA [liu-NeurIPS-2023]. The examples include both relevant explanations as well as wrong explanations, e.g. identifying artifacts on real videos.
@@ 8 Additional Results
In Figure 11, we showcase some examples of how the ViT-based click predictor compares with the ground-truth click locations. We observe that the predictor is able to precisely locate visual artifacts.
Figure 11: Qualitative examples of our click predictor. The green dots represent the ground-truth locations and the blue dots represent the corresponding predictions. Best viewed in color.
@@ 9 Training Environments
We have worked on multiple environments for our experiments. For the in-context learning experiments, we used a Tesla V100-SXM2 GPU with 32GB VRAM. Phi-3-Vision [abdin-arxiv-2024] and LLaVA [liu-NeurIPS-2023] were fine-tuned on a single H100 GPU with 80GB VRAM. BLIP-2 [li-ICML-2023] was fine-tuned using an RTX 4090 GPU with 24GB VRAM. The same training environment as for BLIP-2 was used for the click predictors.